OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -3 - 2 cos θ The answer I chose is y-axis only. But the other answers are no symmetry, origin only and x-axis only.

OpenStudy (anonymous):

Am I right?

OpenStudy (bohotness):

i think so

OpenStudy (anonymous):

@Helpme323 @iGreen @mathmath333 help?

OpenStudy (anonymous):

Or is it x-axis? I think it might be that one after looking back at the graph.

OpenStudy (anonymous):

@aum @CuteGirl903k @IM5er101 @just_one_last_goodbye @No.name @puppylover123 @TheSmartOne I just need someone to check this. I think it might be the x-axis. Am I right?

OpenStudy (anonymous):

sorry idk this!! : /

OpenStudy (anonymous):

\[\begin{array}{c|l} r(\theta)=\cdots&\text{Symmetry}\\ \hline r(-\theta)&\text{x axis}\\ \hline r(\pi-\theta)&\text{y axis}\\ \hline -r(\theta)&\text{origin} \end{array}\] Given \(r=-3-2\cos\theta\), you would check to see if this expression is equivalent to any of the other forms. \[r(\theta)=-3-2\cos\theta\\ r(-\theta)=-3-2\cos(-\theta)=\cdots\] If these are the same, you have symmetry about the x axis.

OpenStudy (anonymous):

It's okay @puppylover123 thanks anyway

OpenStudy (anonymous):

Im good at english tho!! lol

OpenStudy (anonymous):

if u have a question in english post it and tag me :)

OpenStudy (anonymous):

Thank you @SithsAndGiggles !! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@puppylover123 Thanks I'll keep that in mind!